OpenStudy (anonymous):

Jacob walked for x miles and then cycled for y miles between his home and the library. The number of miles Jacob cycled was 3 miles more than twice the number of miles he walked. The total number of miles Jacob walked and cycled was 6 miles. Which graph best shows the number of miles Jacob walked and cycled?

OpenStudy (anonymous):

i believe its y = 2x + 3 x + y = 6 substitute first equation into second equation: x + (2x + 3) = 6 3x + 3 = 6 3x = 3 x = 1 using first equation: y = 2(1) + 3 = 5 check using second equation: 1 + 5 = 6 checks x = 1, y = 5 or (1, 5) intersection point Graph (4) is the only one that intersects at (1, 5). but i want to make sure

imqwerty (imqwerty):

yes correct :)

OpenStudy (anonymous):

yay . tyvm

imqwerty (imqwerty):

np (: